OpenStudy (anonymous):

prove that (2 - 2cos^2 x -sin x) / (2sin x cos x - cos x) =tan x I am a bit stumped by this

OpenStudy (anonymous):

(2 - 2cos^2 x -sin x) / (2sin x cos x - cos x) =(2 (1- cos^2 x)-sin x) / (2sin x cos x - cos x) =(2 sin^2x -sinx)/ (2sin x cos x - cos x) =sinx(2sinx-1)/(cosx(2sinx -1)) =sinx/cosx =tanx

OpenStudy (anonymous):

Thank you for your help